# L4P: Low-Level 4D Vision Perception Unified

### v1

@@ Abstract
The spatio-temporal relationship between the pixels of a video carries critical information for low-level 4D perception. A single model that reasons about it should be able to solve several such tasks well. Yet, most state-of-the-art methods rely on architectures specialized for the task at hand. We present L4P (pronounced “LAP”), a feedforward, general-purpose architecture that solves low-level 4D perception tasks in a unified framework. L4P combines a ViT-based backbone with per-task heads that are lightweight and therefore do not require extensive training. Despite its general and feedforward formulation, our method matches or surpasses the performance of existing specialized methods on both dense tasks, such as depth or optical flow estimation, and sparse tasks, such as 2D/3D tracking. Moreover, it solves all those tasks at once in a time comparable to that of individual single-task methods.
@@ 1 Introduction
Large collections of videos are our most complete and compact source of priors about the world. Much like text did for large-language models, the corpus of videos we amassed over the years allowed video-language models (VLMs) to produce remarkable zero-shot results on high-level vision tasks such as video captioning, video question answering, and others. However, zero-shot, low-level 3D and 4D vision perception tasks, such as depth from video, tracking, optical flow, and others remain a challenge. Pretrained video diffusion models fine-tuned on target-domain data showed potential on dense vision perception tasks (e.g., depth [27, 20], flow [43], etc.), but the fine-tuning makes them task-specific, and therefore limits their ability to leverage priors across multiple tasks at once. Sparse vision perception tasks, such as tracking, are even more challenging to tackle with a general foundation model, because their representation does not fit naturally into data structures like dense pixel planes. As a result, they have been historically addressed with carefully designed approaches, which are often optimization-based [59], or with specialized architectures [26, 8].
@@
Can we leverage the priors learned from a large body of video data and solve multiple low-level 4D vision perception tasks, both dense and sparse, with a unified architecture, in a zero-shot way?
@@
This goal presents multiple challenges. To effectively learn and share priors across tasks—including those beyond 4D perception—we need an architecture with a strong backbone shared across tasks. This architecture should also be general to allow for pretraining on auxiliary tasks. Not least, dense and sparse tasks require fundamentally different representations, e.g., dense 2D planes vs. 3D tracks.
@@
Figure 1: We propose L4P, a single, general-purpose architecture that solves several low-level 4D perception tasks with zero-shot generalization capabilities. We show that a pre-trained video encoder can be combined with lightweight heads and surpass or at least match the performance of SOTA methods, which use specialized architectures and are trained to solve individual tasks.
@@
We fulfill these requirements by combining a pretrained video masked auto-encoder (VideoMAE) [56, 58] with per-task, lightweight heads (Figure 2). VideoMAEs have been successfully employed for a variety of mid- and high-level vision tasks [58], but their ability to capture the spatio-temporal relationship between pixels is underexplored in the context of low-level 4D perception. We choose a VideoMAE as the backbone for our system because of the powerful priors it learned in its pretraining. Moreover, VideoMAEs offer a feedforward, online mechanism to tokenize videos within a relatively small computational envelope. For dense tasks, we couple the VideoMAE with heads based on the dense prediction transformer (DPT), which has been shown to perform well on depth estimation, image segmentation, and others [41]. For sparse tasks we focus on tracking, and specifically, on the track-any-point (TAP) formulation [7]. We posit that tracking is important for perception because understanding fine-grained, complex motions and the physical interaction between objects is critical to downstream applications, including 3D reconstruction [32, 57] and robotics manipulation [70, 66, 67, 1]. We formulate the problem of estimating tracks as that of predicting, for queried pixels, 2D heatmaps with associated depth and visibility tokens. This is the mechanism that allows us to tackle sparse and dense tasks within a unified framework.
@@
Our formulation presents several desirable properties and advantages. First, the pretrained VideoMAE model allows us to tap into priors learned from large datasets—potentially different and more varied than those typically used for low-level 4D perception. It also affords us efficient computation: our system solves all tasks in roughly 300ms for a 16-frame video chunk ($\sim$19ms/frame), which is comparable to, or faster than, methods specialized for each task (see Table 1 in the Supplementary). Moreover, combining it with per-task heads allows us to train a relatively small number of parameters for new tasks, which we show by freezing the system and adding a head for motion-based segmentation. Lastly, but perhaps most importantly, breaking the architecture into a general VideoMAE and per-task heads offers a mechanism to solve both dense and sparse tasks with a unified model (Figure 1). Despite being general, our architecture performs on par with, or better than, state-of-the-art methods. This is remarkable because the baselines we compare against are task-specific, carefully designed for, and specialized to excel at their respective tasks. Finally, VideoMAEs are already used as encoders for VLMs [64, 34], and we speculate that training them to reason about low-level 4D perception may impart those capabilities to the downstream VLMs they may be used with, though validating this statement is outside the scope of our paper.
@@
Figure 2: We split our architecture into a powerful encoder, which extracts tokens from the input video, and per-task lightweight heads. For sparse tasks, such as 3D tracking, we define additional query tokens, namely the point we wish to track, $\mathcal{P}$, with the corresponding feature token, $\mathcal{F}$, and output tokens (heatmap $\mathcal{H}$, depth $\mathcal{D}$, and visibility $\mathcal{V}$). Figure 3 further details the processing performed by the sparse task head.
@@ 2.1 Foundation models for vision perception
Self-supervised pre-training of large models on huge unlabeled data has shown great success on vision tasks. Among them, Vision Transformers (ViT) [29] pretrained with masked autoencoding (MAE) [19] have become a major choice to fine-tune for many vision tasks, SegmentAnything (SAM) [28] being a notable example. VideoMAE [56] and MAE-ST [12] adopt MAE on videos for spatio-temporal representation learning. VideoMAEv2 [58] introduces a dual masking strategy that allows them to efficiently scale up the model to a billion parameters and to effectively leverage priors from large data. Their representation shows strong performance on action recognition, and more recently is combined with large language models to enable multimodal video understanding [64, 34]. However, their uses remain unexplored for the low-level 4D perception tasks, which we address in this work.
@@ 2.2 Dense Prediction Tasks
We review related works for the dense prediction tasks of depth, flow and motion-based segmentation.
Depth Estimation. Depth perception is critical to many applications, and depending on the use cases many solutions exists for this task. Traditional stereo approaches [14, 15, 46], as well as their more recent deep learning counterparts [75, 21, 44], take two or more posed input views and leverage the underlying 3D correspondences, which limit their uses to only static scenes. Depth estimation from a single image [11, 40, 41, 73, 74] has made impressive progress, but still suffer from 3D and temporal inconsistency, limiting their utility for video settings. In this work, we focus on video depth estimation, which, unlike multi-view stereo (MVS), focuses on dynamic scenes and assumes no provided camera poses. Some early works rely on test-time optimization [35, 30] to enforce consistency. More recent works introduce feedforward [52, 63] and diffusion-based [48, 20, 76] solutions.
Since depth annotations are limited, many approaches exploiting learned priors from foundation models. DINOv2 [38] pretraining allows them to achieve competitive single-image depth estimation quality with only a lightweight head and minimal fine-tuning. DUSt3R [60] and MASt3R [33] show that cross-view completion pre-training [65] can provide strong prior for cross-view reasoning. Diffusion models are another powerful source of priors, where Marigold [27] and DepthCrafter [20] demonstrate impressive detailed outputs in single-image and video depth estimation respectively. Like in previous works, we adopt large-scale pretraining by using a video encoder pre-trained with MAE; however, we solve this task along with other 4D perception tasks in a unified framework.
Optical flow estimation. Though straightforward when posed as a dense prediction task, optical flow traditionally requires specialized architectures leveraging domain knowledge to be competitive [53, 50, 61, 69]. Similar to depth estimation, recent works adopt powerful priors from large-scale pretraining, e.g. via diffusion priors [43] or large-scale cross-view completion pretraining [65]. Most related to ours are approaches that take multi-frame inputs [42, 23, 49, 10]. They generally showcase improved accuracy and stability in predictions by leveraging temporal coherence; however they all rely on heavily customized networks, in contrast to ours using only generic architectures.
Motion-based segmentation. Early learning-based approaches rely on combining appearance features and optical flow to solve this task [1, 13, 22, 55, 54]. Recent works extract geometric properties from optical flow before using it to train a binary classification network. For example, MoA-Net [4] compensates for the camera rotation in the optical flow, and RigidMask [72] goes further by extracting optical expansion [71] signal from flow and using single-image depth priors. However, they are prone to noises in the estimated flow and are limited by their two-frame formulation. We show that using priors from our video-based approach, we can achieve good performance by only training a task-specific head for this task.
@@ 2.3 Sparse Prediction Tasks
Tracking Any Point (TAP) in a video for long durations is fundamental for understanding complex dynamics of the world and has many applications [67, 1, 32, 57, 6]. Particle Video [18], PIPs [3] and TAP-Net [7] lay the initial foundation by adopting several ideas from optical flow approaches like building cost-maps between query points and image features, iterative estimation of tracks, etc. On the other hand, OmniMotion [59] optimizes a volumetric representation for each video to solve this task, but its time-consuming nature limits its applicability. TAPIR [8] introduces the idea of coarse-to-fine track estimation with a global temporal refinement strategy and BootsTAPIR [9] further improves it by adopting a self-supervised learning approach. CoTracker [26] proposes to jointly track multiple points to leverage spatial correlations and achieve strong 2D tracking performance while still being an online approach. SpaTracker [68] introduces one of the first feedforward 3D point tracking approaches by using depthmaps from a depth estimator to uplift pixels to 3D and leveraging CoTracker’s tracking formulation to track points in 3D.
Unlike previous approaches that rely on carefully designed, specialized architectures, our solution solves both the 2D and 3D point tracking tasks in a single unified framework shared with other tasks. We share some motivations with Tracking at Any Granularity [17], a concurrent work. While they focus on training a large model from scratch on a large collection of datasets for tracking various types of queries, we focus on leveraging large-scale pre-trained models to solve the point tracking task as one of the many capabilities afforded by our unified framework.
@@ 3 Method
Figure 3: Detailed structure for the sparse head. The tokens extracted from the video and the I/O tokens (query $\mathcal{P}$oint token, query point $\mathcal{F}$eature token, and query point $\mathcal{H}$eat map, $\mathcal{D}$epth, and $\mathcal{V}$isibility) tokens are processed by a SAM-style two-way attention layer. The outputs are then reshaped and resized. The resulting per-frame featuremaps and $\mathcal{H}$eat map, $\mathcal{D}$epth, and $\mathcal{V}$isibility tokens are combined via a dot-product. We also implement a memory mechanism that combines the video tokens and the $\mathcal{F}$eature token from overlapping windows.
We provide an overview of our approach in Figure 2. Our model uses a pre-trained ViT-based video encoder [58] (Section 3.1) to capture spatio-temporal features in an RGB video clip of length $T$. We use lightweight task-specific heads that decode the video features for low-level 3D/4D perception tasks. For pixel-wise dense tasks like depth, flow, and motion-based segmentation, we design a DPT-based head [41], originally only designed for images, to work with videos (Section 3.2). For the sparse task of tracking any pixel in a video, we take inspiration from the head architecture proposed in SAM [28] (Section 3.3). Given a pixel queried in any frame of the input video, we extend the head, also originally designed to work for images, to decode video tokens into a 2D trajectory, its depth with respect to the camera, and the track visibility in each frame (Section 3.3.1). The VideoMAE architecture extracts video tokens from windows of fixed length $T$, and cannot process longer sequences. To allow for an online approach, we propose a memory mechanism to track points in arbitrarily long videos (Section 3.3.2).
@@ 3.1 Video Masked Auto-Encoders
Motivated by scalable and powerful pre-training methods and architectures, we use the ViT-based video encoder from VideoMAEv2 [58], which was pre-trained using the masked auto-encoding task. The encoder works with videos of size $T\times H\times W$, uses a spatio-temporal patch-size of $t\times h\times w$ and cube embedding [56] to transform an input video into a sequence of tokens. These are then processed by the ViT architecture with spatio-temporal attention to generate the video tokens $\mathcal{S}\in\mathbb{R}^{P\times C}$, where $P$ is the number of tokens and $C$ is the embedding dimension. We run the video encoder only once per video clip. Once they are encoded, we can apply the lightweight heads to decode the tokens to the desired output. For the point tracking task, we can independently prompt these tokens to track many points in parallel.
@@ 3.2 Dense Prediction Heads
Dense prediction tasks produce outputs with spatial dimensions aligned with their inputs, typically at the same resolution $H\times W$. A wide array of common computer vision problems can be formulated as dense prediction tasks. In this work, we explore depth estimation, optical flow estimation, and motion-based segmentation as examples.
It is often critical to capture both local and global spatial structures for dense prediction tasks to succeed. We opt to adapt DPT [41] as our dense prediction head due to its proven performance and efficiency on single-image depth estimation. DPT progressively assembles and combines tokens from various layers inside transformers to produce full-resolution predictions. To leverage the 3D tokens from VideoMAE and to enable temporal reasoning, we replace all 2D convolutions inside the DPT head with 3D convolutions. We find that this modification is enough to bring in temporal consistency with minimal computation overhead. The DPT heads for each of the dense tasks differ only in the final layer, which outputs one channel for depth and motion-based segmentation, and two channels for optical flow.
For videos longer than $T$ frames, we run inference with stride $T/2$. To enforce consistency, we use an affine transformation to align the depth predictions for frames at the overlap of consecutive windows. This strategy has no effect on the individual windows for relative depth, but it greatly improves long-term temporal consistency. For optical flow and motion-based segmentation, we simply overwrite the overlapping predictions.
Figure 4: We show results for all perception tasks supported by L4P. Both examples show dynamic scenes with camera and object motions.

### v2

@@ Abstract
The spatio-temporal relationship between the pixels of a video carries critical information for low-level 4D perception tasks. A single model that reasons about it should be able to solve several such tasks well. Yet, most state-of-the-art methods rely on architectures specialized for the task at hand. We present L4P, a feedforward, general-purpose architecture that solves low-level 4D perception tasks in a unified framework. L4P leverages a pre-trained ViT-based video encoder and combines it with per-task heads that are lightweight and therefore do not require extensive training. Despite its general and feedforward formulation, our method matches or surpasses the performance of existing specialized methods on both dense tasks, such as depth or optical flow estimation, and sparse tasks, such as 2D/3D tracking. Moreover, it solves all tasks at once in a time comparable to that of single-task methods.
@@ 1 Introduction
Large collections of videos are our most complete and compact source of priors about the world. Much like text did for large-language models, the corpus of videos we amassed over the years allowed video-language models (VLMs) to produce remarkable zero-shot results on high-level vision tasks such as video captioning, video question answering, and others. However, zero-shot, low-level 3D and 4D vision perception tasks, such as depth from video, tracking, optical flow, and others remain a challenge. Pre-trained diffusion models fine-tuned on target-domain data showed potential on dense vision perception tasks (e.g., depth [28, 21], flow [44], etc.), but the fine-tuning makes them task-specific, and therefore limits their ability to leverage priors across multiple tasks at once. Sparse vision perception tasks, such as tracking, are even more challenging to tackle with a general foundation model, because their representation does not fit naturally into data structures like dense pixel planes. As a result, they have been historically addressed with carefully designed approaches, which are often optimization-based [63], or with specialized architectures [27, 8].
@@
Can we leverage the priors learned from a large body of video data and solve multiple low-level 4D vision perception tasks, both dense and sparse, with a unified architecture with strong generalization abilities?
@@
This goal presents multiple challenges. To learn and share priors across tasks—including those beyond 4D perception—we need a strong backbone shared across tasks. The architecture should also be general to allow for pre-training on auxiliary tasks. Not least, dense and sparse tasks require fundamentally different representations, e.g., dense 2D planes vs. 3D tracks.
@@
Figure 1: We propose L4P, a general-purpose architecture that solves several low-level 4D perception tasks. Building on a pre-trained video encoder and lightweight per-task heads, our unified model matches or surpasses the performance of SOTA methods specialized to solve individual tasks. L4P can easily be extended to additional tasks, (e) and (f), without compromising performance.
@@
We tackle these challenges by combining a pre-trained video masked auto-encoder (VideoMAE) [59, 62] with per-task, lightweight heads (Figure 2). VideoMAEs have been successfully employed for a variety of mid- and high-level vision tasks [62], but their ability to capture the spatio-temporal relationship between pixels is underexplored in the context of low-level 4D perception. VideoMAEs offer a feedforward, online mechanism to tokenize videos within a small computational envelope. For dense tasks, we couple VideoMAE with heads based on the dense prediction transformer (DPT), which has been shown to perform well on depth estimation, image segmentation, and others [42]. For sparse tasks we focus on tracking, and specifically, on the track-any-point (TAP) formulation [7]. We posit that tracking is important for perception because understanding fine-grained, complex motions and the physical interaction between objects is critical to downstream applications, including 3D reconstruction [33, 61] and robotics manipulation [76, 72, 73, 1]. However, the point tracking tasks have historically been addressed using specialized architectures [27, 8, 74]. To unify them with other dense tasks, we formulate the problem of estimating tracks as that of predicting, for queried pixels, 2D heatmaps with associated depth and visibility tokens. This is the mechanism that allows us to tackle sparse and dense tasks within a unified model. To implement it, we augment a general-purpose head with a novel memory mechanism to track points for arbitrarily long videos. Thanks to the training curriculum we propose, and to VideoMAE’s priors, our unified model finetuned for multiple tasks at once on a small collection of synthetic datasets shows strong generalization to real data.
@@
Our formulation presents several desirable properties and advantages. First, the pre-trained VideoMAE model allows us to tap into priors learned from large datasets that are potentially different and more varied than those typically used for low-level 4D perception. It also affords us efficient computation: our system solves all tasks in roughly 300ms for a 16-frame video chunk ($\sim$19ms/frame), which is comparable to, or faster than, methods specialized for each task (see Table 1 in the Supplementary). Moreover, combining it with per-task heads allows us to train a relatively small number of parameters for new tasks, which we show by freezing the system and adding a head for motion-based segmentation and camera pose estimation (marked in green in Figures 1 and 2). Lastly, but perhaps most importantly, breaking the architecture into a general VideoMAE and per-task heads offers a mechanism to solve both dense and sparse tasks with a unified model (Figure 1). Despite being general, our architecture performs on par with, or better than, state-of-the-art methods. This is remarkable because the baselines we compare against are task-specific, carefully designed for, and specialized to excel at their respective tasks. Finally, VideoMAEs are already used as encoders for VLMs [70, 34], and we speculate that training them to reason about low-level 4D perception may impart those capabilities to the downstream VLMs they may be used with, though validating this statement is outside the scope of our paper.
@@
Figure 2: We use a pre-trained video encoder as tokenizer, and combine it with per-task lightweight heads. For sparse tasks we define additional query tokens: the point to track, $\mathcal{P}$, the corresponding feature token, $\mathcal{F}$, and output tokens (heatmap $\mathcal{H}$, depth $\mathcal{D}$, and visibility $\mathcal{V}$).
@@ 2.1 Foundation models for vision perception
Self-supervised pre-training of large models on huge unlabeled data has shown great success. Among them, Vision Transformers (ViT) [30] pre-trained with masked autoencoding (MAE) [20] have become a common choice to fine-tune for many vision tasks, Segment Anything (SAM) [29] being a notable example. VideoMAE [59] and MAE-ST [12] adopt MAE on videos for spatio-temporal representation learning. VideoMAEv2 [62] introduces a dual masking strategy that allows them to efficiently scale up the model to a billion parameters. Their representation is effective for action recognition, and more recently is combined with language models [70, 34]. However, their uses remain unexplored for the low-level 4D perception tasks, which we address in this work.
@@ 2.2 Dense Prediction Tasks
Depth Estimation. Stereo approaches [14, 15, 47, 80, 22, 45] estimate depth by estimating correspondences between posed images, which limit their uses to static scenes. Depth estimation from a single image [11, 41, 42, 78, 79] has made significant progress, but suffers from 3D and temporal inconsistency in a video setting. We focus on estimating depth from videos of dynamic scenes, for which several solutions exist [35, 31, 54, 69]. To tackle the data scarcity problem, the state-of-the-art approaches leverage priors from foundation models: DepthAnything [78, 79] uses DINOv2 [38], DUSt3R [65] uses cross-view completion pre-training [71], and Marigold [28] and DepthCrafter [21] use image and video diffusion priors respectively. We benefit from large-scale pre-training by using a video encoder pre-trained with MAE.
Optical flow estimation. Though straightforward when posed as a dense task, optical flow traditionally requires specialized architectures [55, 52, 66, 75]. Most related to ours are approaches that take multi-frame inputs [43, 24, 50, 10]. Similar to depth estimation, recent works adopt priors from large-scale pre-training, e.g. via diffusion priors [44] or cross-view completion pre-training [71].
Motion-based segmentation. While early learning-based approaches rely on combining appearance features with flow [1, 13, 23, 58, 57], more recent works [4, 77] extract geometric properties from flow before using it to train a classification network. However, they are affected by noisy flow and are limited by their two-frame formulation.
Camera pose estimation. Classical structure-from-motion solutions [46, 51] work robustly and accurately in constrained settings, while learning-based solutions [67, 56, 65, 60, 81, 64] address challenging scenarios like limited image observations, dynamic and texture-less regions, etc. We cast camera pose estimation as a dense task by representing cameras as dense bundles of rays [17, 82] using 6-D Plücker coordinates [40] and estimating them using our dense head.
We treat motion-based segmentation and camera pose estimation as additional tasks, by leveraging the priors from our video model trained on depth, flow and tracking tasks, and only training task-specific heads. Thus, we maintain main tasks performance while supporting additional tasks.
@@ 2.3 Sparse Prediction Tasks
Tracking Any Point (TAP) in a video has many applications [73, 1, 33, 61, 6]. Particle Video [19], PIPs [3] and TAP-Net [7] lay the initial foundation by adopting ideas from optical flow approaches. On the other hand, OmniMotion [63] optimizes a volumetric representation for each video to solve this task and is time-consuming. TAPIR [8] introduces the idea of coarse-to-fine track estimation, BootsTAPIR [9] further improves it by adopting a self-supervised learning, and CoTracker [27] proposes to jointly track multiple points to leverage spatial correlations. SpaTracker [74] is one of the first feedforward 3D point tracking approaches. It uses estimated depth maps to uplift pixels to 3D and leverages CoTracker’s tracking formulation to track points in 3D.
Unlike previous approaches with specialized architectures, our solution solves point tracking in a unified way with other tasks. We share some motivations with Tracking at Any Granularity [18], a concurrent work. While they focus on training a large model from scratch on a large collection of datasets for tracking various types of queries, we choose to leverage a large-scale pre-trained model that lends itself to a wide range of tasks.
@@ 3 Method
Figure 3: Sparse head. The video and I/O tokens (query $\mathcal{P}$oint, query point $\mathcal{F}$eature, and query point $\mathcal{H}$eat map, $\mathcal{D}$epth, and $\mathcal{V}$isibility) are processed by a two-way attention layer. The reshaped and resized per-frame feature maps and processed output tokens are combined via a dot product. We also introduce a memory mechanism to combine video and the $\mathcal{F}$eature tokens from time $t$ and $t+1$.
We provide an overview of our approach in Figure 2. Our model uses a pre-trained ViT-based video encoder [62] (Section 3.1) to capture spatio-temporal features in an RGB video clip of length $T$. We use lightweight task-specific heads that decode the video features for low-level 3D/4D perception tasks. For pixel-wise dense tasks like depth, flow, motion-based segmentation, and camera rays estimation (for camera pose estimation), we propose an extension of the DPT architecture [42] that allows us to use them for videos, instead of just images, as in their existing formulation (Section 3.2). For the sparse task of tracking any pixel in a video, we take inspiration from the head architecture proposed in SAM [29] (Section 3.3). Given a pixel queried in any frame of the input video, we extend the head, also originally designed to work for images, to decode video tokens into a 2D trajectory, its depth with respect to the camera, and its visibility in each frame (Section 3.3.1). VideoMAE extracts video tokens from temporal windows of fixed length $T$, and cannot process longer sequences. We propose a memory mechanism to track points in arbitrarily long videos (Section 3.3.2).
@@ 3.1 Video Masked Auto-Encoders
Motivated by its scalable and powerful pre-training and architecture, we use the ViT-based video encoder from VideoMAEv2 [62]. The encoder works with videos of size $T\times H\times W$ and uses a spatio-temporal patch size of $t\times h\times w$. It uses cube embedding [59] to transform an input video into a sequence of tokens, which are then processed by the spatio-temporal attention blocks to generate video tokens $\mathcal{S}\in\mathbb{R}^{P\times C}$, where $P$ is the number of tokens and $C$ is the embedding dimension. We run the video encoder only once per video clip, and then apply the lightweight heads to decode the tokens to the desired outputs. For the point tracking task, we can independently prompt these tokens to track many points in parallel.
@@ 3.2 Dense Prediction Heads
Dense prediction tasks produce outputs with spatial dimensions aligned with their inputs, typically at the same resolution $H\times W$. A wide array of common computer vision problems can be formulated as dense prediction tasks. In this work, we explore depth estimation, optical flow estimation, motion-based segmentation, and camera pose estimation as examples.
We adopt the DPT [42] architecture for dense prediction due to its proven performance and efficiency on single-image depth estimation. DPT progressively assembles and combines tokens from various layers, thus capturing both local and global spatial structures. We adapt DPT to work with the 3D tokens from VideoMAE by mapping the video tokens to 3D feature maps and enabling temporal reasoning by introducing 3D convolutions inside the DPT head. We find that this modification is enough to bring in temporal consistency with minimal computation overhead. For most of our dense tasks, the DPT heads differ only in the final layer, which outputs one channel for depth and motion-based segmentation, and two channels for optical flow. We cast camera pose estimation as a dense task by estimating a bundle of rays for each camera, and recover camera parameters from the estimated rays [82] (see Supplementary for more details).
For videos longer than $T$ frames, we run inference with stride $T/2$ and enforce consistency in the overlapping frames between temporal windows. For depth, we align the predictions with an affine transformation. This strategy has no effect on the individual windows for relative depth, but can greatly improve long-term temporal consistency. For optical flow and motion-based segmentation, we simply overwrite the predictions in the overlap between windows. For pose estimation, we align the camera trajectories in the overlapping windows with a similarity transformation.
Figure 4: We show results for perception tasks for scenes with both camera and object motion.